Drawing a heart and a diamond in any order in 2 consecutive draws when the first card is returned to the deck and the deck is shuffled after the first draw.
My Answer: 26/52*13/52=338/2704=1/8

Answer by rfer(16322)

(13/52)(13/52)=169/2704=1/16
Your first draw has a chance of 13/52 and your second draw has a chance of 13/52 since the first card was replaced in the deck.